OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical 1a. (2, -5) and (-1, -1) 1b. (7, 1) and (7, 2) 1c. (1, 7) and (-2, -2) 1d. (-1, -5) and (-4, -5)

OpenStudy (anonymous):

Do you know the slope formula?

OpenStudy (anonymous):

i dont

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Ok, very quick and simple: Given two points: \((x_1,y_1)\) and \((x_2,y_2)\) The slope of the line between them is found using the formula: \[Slope = \frac{y_2-y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

So take the points you have there, and plug them in. That will give you the slope between them.

OpenStudy (anonymous):

We can check your answers as you find them

OpenStudy (anonymous):

I think I got all the answers, but I'm having trouble figuring out whether the slopes are horizontal or vertical

OpenStudy (anonymous):

If the slope comes out to be \(\frac{0}{k}\) where k is some other number then the slope is horizontal. If the slope comes out to he \(\frac{k}{0}\) where k is some other number then the slope is vertical.